OpenStudy (adi3):

Will medal Please help for f(x)^3 find the simplest form: (i) F(x) + 2

OpenStudy (adi3):

@iambatman

OpenStudy (adi3):

@imqwerty

OpenStudy (adi3):

@phi

OpenStudy (phi):

there is not enough info. There must be more to the question.

OpenStudy (adi3):

Please help for f(x)^3 find in simplest form: (i) F(x) + 2

OpenStudy (anonymous):

what is your f(x), what are the other options?

OpenStudy (phi):

you should post a screen shot of the actual question.

OpenStudy (adi3):

fine i will post a scrren shot

OpenStudy (adi3):

wait

OpenStudy (adi3):

is this clear enough

OpenStudy (adi3):

the first question

OpenStudy (anonymous):

You should start off by plotting f(x)=x^3 first

OpenStudy (adi3):

then

OpenStudy (phi):

ok, they are saying if you have the curve \[ f(x)= x^3\] write the equation of a new curve, that we get by adding 2: \[ F(x) = x^3 + 2\] and then they want you to graph both curves.

OpenStudy (anonymous):

Then depending on what you add or subtract from f(x), we can accordingly shift our graph

OpenStudy (phi):

Do you know how to graph \( y=x^3 \)?

OpenStudy (adi3):

yes

OpenStudy (adi3):

and @Nishant_Garg how did u do tht

OpenStudy (anonymous):

graph of x^3 will look something like